### v1

@@ 5 Discussion and conclusions
In this paper thermal evolution of a newborn strange star within the bag model of non-interacting quark matter is considered, taking into account kinetics of pairs created in its electrosphere. Numerical solution of the heat transfer equation in neutrino transparent regime shows that the surface temperature of newborn strange star reduces from initial $10^{11}$ K to $10^{10}$ K on a very short timescale $\Delta t\sim 10^{-7}$ seconds. At the same time, interior temperature starts to change on longer timescale, about $10^{-3}$ seconds, due to the URCA process. The reason of high temperature gradient, established near the surface is a combination of high power of the the Schwinger process and low thermal conductivity of free quarks. Hence the surface of newborn strange star cannot sustain high luminosity on long timescales, as required for the explanation of extreme astrophysical transient events, such as gamma-ray bursts.
@@
Here we did not consider subsequent evolution of pair outflow at large distances from the surface of the star, as this topic is already covered in the literature by Aksenov et al. (2004, 2005). Note that the luminosity in pairs decreases from $10^{40}$ erg/s at $t\sim 1$ second to $10^{36}$ erg/s at $t\sim 10^{4}$ seconds. This is precisely the luminosity range where a transition from black body-like spectrum (from thermalization of pairs due to large optical depth of the outflow) to annihilation line occurs (Aksenov et al., 2004). This effect is potentially observable and may represent the unique feature of quark star cooling.
@@
We also considered the effects of neutrino trapping at high temperatures and found that simplified integral models, such as the ones presented by Berdermann et al. (2004); Li et al. (2021) do not account for the sharp temperature gradient at the surface of quark star. They fail to estimate the surface temperature and therefore cannot be used to obtain pair luminosity as a function of time. To our knowledge, detailed calculations of the dynamics of neutrinosphere inside a newborn strange star is not performed yet. We are developing a spherical Boltzmann code for neutrino transfer ((Vereshchagin & Aksenov, 2017)) in order to close this gap. Nevertheless, our toy model for neutrino transparent shell, described in section above, clearly shows that the conclusion obtained for the neutrino transparent case holds.
@@
Note, that a true QCD phase of strange quark matter is still uncertain. If color superconductivity takes place, it can increase thermal conductivity of quarks (Alford et al., 2008). Thermal conductivity in the two-color–flavor superconductor (2SC) phase is similar to that of non-interacting quarks. The most interesting case is color-flavor-locked (CFL) phase, where thermal conductivity is much larger. Strange star in CFL phase does not contain electrons, but the surface electron layer appears, because of the number of massive quarks (s quarks) is reduced near the boundary relative to the number of massless quarks. In this case the structure of electrosphere differs significantly (Usov, 2004). We are planning to treat quark stars in CFL phase in the second paper of this series.
@@
Acknowledgements. This work is supported within the joint BRFFR-ICRANet-2025 funding programme under the grant No. F25ICR-001.

### v2

@@ V Discussion and conclusions
In this paper thermal evolution of a newborn strange star made of unpaired strange quark matter is considered, taking into account kinetics of pairs created in its electrosphere. Numerical solution of the heat transfer equation in neutrino transparent regime shows that the surface temperature of newborn strange star reduces from initial $10^{11}$ K to $10^{10}$ K on a very short timescale $\Delta t\sim 0.1$ seconds. At the same time, interior temperature starts to change on longer timescale, about $10^{-3}$ seconds, due to the URCA process. The reason of high temperature gradient, established near the surface is a combination of high power of the the Schwinger process and low thermal conductivity of unpaired quarks. Hence the surface of newborn strange star cannot sustain high luminosity on long timescales, as required for the explanation of extreme astrophysical transient events, such as gamma-ray bursts.
@@
Here we did not consider subsequent evolution of pair outflow at large distances from the surface of the star, as this topic is already covered in the literature 2004ApJ…609..363A ; 2005ApJ…632..567A . Note that the luminosity in pairs decreases from $10^{46}$ erg/s at $t\sim 1$ second to $10^{36}$ erg/s at $t\sim 10^{5}$ seconds. This is precisely the luminosity range where a transition from black body-like spectrum (from thermalization of pairs due to large optical depth of the outflow) to annihilation line occurs 2004ApJ…609..363A . This effect is potentially observable and may represent the unique feature of quark star cooling.
@@
We also considered the effects of neutrino trapping at high temperatures and found that simplified integral models, such as the ones presented bin2004csqn.conf..399B ; 2021ApJ…922..214L do not account for the sharp temperature gradient at the surface of quark star. They fail to estimate the surface temperature and therefore cannot be used to obtain pair luminosity as a function of time. To our knowledge, detailed calculations of the dynamics of neutrinosphere inside a newborn strange star is not performed yet. We are developing a spherical Boltzmann code for neutrino transfer (2017rkt..book…..V ) in order to close this gap. Nevertheless, our toy model for neutrino transparent shell, described in section above, clearly shows that the conclusion obtained for the neutrino transparent case holds.
Note, that a true QCD phase of strange quark matter is still uncertain. If color superconductivity takes place, it can increase thermal conductivity of quarks RevModPhys.80.1455 .
Thermal conductivity in the two-color–flavor superconductor (2SC) phase was revised in PhysRevC.90.055205 . It was shown that at high temperatures quark-quark interactions do not contribute to the total thermal conductivity and electron-quark interactions become the main ones. The corresponding thermal conductivity $\kappa_{e}$ for $T=10^{11}$ K and $\mu_{q}=300$ MeV is of the order of $10^{21}$ erg/s/cm/K. Comparing the last value with the unpaired case studied in this paper $\kappa\simeq 10^{22}$ erg/s/cm/K we see that the strange star surface made of 2SC quarks can not preserve high temperature for a long time.
@@
According to PhysRevC.66.015802 thermal conductivity of the color-flavor-locked (CFL) phase can be of the order of $10^{33}$ erg/s/cm/K. Strange star in CFL phase does not contain electrons, but the surface electron layer can appear, because of the number of massive quarks (s quarks) is reduced near the boundary relative to the number of massless quarks. In this case the structure of electrosphere differs significantly PhysRevD.70.067301 .
We will consider strange quark stars with color superconductivity in the next paper.
@@
Acknowledgements. We thank the anonymous referee for his comments which improved the paper. This work is supported within the joint BRFFR-ICRANet-2025 funding programme under the grant No. F25ICR-001.